OpenStudy (anonymous):

help write as fraction? 1.) 0.34 2.) 0.24 3.) 0.45 4.) 0.010 5.) 0.2

jimthompson5910 (jim_thompson5910):

0.34 means 34 hundredths so, 0.34 =34/100 = 17/50 make sure you reduce as much as possible